OpenStudy (pottersheep):

How can I make (2-x)^2 into (x-2)^2?

OpenStudy (anonymous):

no need, they are the same thing

OpenStudy (anonymous):

\[(5-2)^2=(2-5)^2\] \[(10-2)^2=(2-10)^2\]

OpenStudy (anonymous):

and \[(x-2)^2=x^2-4x+4=(2-x)^2\]